How do you solve:
What is the maximum amount you can borrow today if it must be repaid in 4 months with simple interest at 8% and you know that at that time you will be able to repay no more than $1500?
Thank you!

Answer by mananth(16949)

Principal = $x
rate = 8%
time = 4 months = 1/3 years
...
Amount to be paid = $1500
..
Amount = Principal + interest
amount = x+0.08*0.33*x
Amount = x+0.0264x
Amount = 1.0264x
...
1.0264x=1500
/1.0264
x=1500/1.0264
=$1461 is what you can borrow
...
m.ananth@hotmail.ca